# A Thorough Assessment of the Non-IID Data Impact in Federated Learning

### v1

@@ VI Design Insights and Opportunities
To the time of writing of this work, no analyses or empirical studies about the effect of the spatiotemporal skew on the performance of FL models exist. Thus, for the first time, we produce experiments to understand how different levels of spatiotemporal non-IIDness affect the prediction power of an FL model. The results show (see Table IX) that high levels of space or time skewness decreases the performance of the models, more specifically when the HD is higher than 0.75 (severe non-IIDness). Thus, researchers may benchmark techniques to deal with space and time skew in FL [40, 11, 49] to determine the behavior under high data heterogeneity levels.
Methods to compare mixed non-IIDness types.
@@
Current tools and methods for synthetic partitioning centralized data into federated data [17, 47, 20, 34, 15] focus on simulating one type of non-IIDness (label, feature, quantity, spatiotemporal skewness). Nevertheless, a more realistic scenario would be combining two or more types of data heterogeneity to evaluate how those mixes can alter the performance of FL models. Therefore, for research in FL purposes, it would be interesting to create methods to partition centralized data into federated clients that permit the control of non-IIDness level for each non-IIDness type at the same time.
VII Conclusions
@@
This study provides a comprehensive empirical analysis of the non-IID effect in FL. We benchmarked five state-of-the-art strategies for handling non-IID data distributions under controlled conditions, exploring label, feature, quantity, and spatiotemporal skewness, with a novel focus on the latter. We aim to standardize the methodology for studying data heterogeneity in FL by using HD to quantify data distribution differences. Our findings reveal the significant impact of label and spatiotemporal skew non-IID types on FL model performance. We also demonstrate that the model’s performance drop appears at a double threshold. When HD is higher than 0.5 and 0.75, a higher damage and a steeper decrease in performance slope occurs. Moreover, our results suggest that the FL performance is heavily affected mainly when the non-IIDness is extreme. Thus, we offer valuable recommendations for researchers to address data heterogeneity. This work represents the most thorough examination of non-IIDness in FL to date, providing a robust foundation for future research in FL.
VIII Acknowledgments
Daniel Mauricio Jimenez G. and Andrea Vitaletti were partially supported by PNRR351 TECHNOPOLE – NEXT GEN EU Roma Technopole – Digital Transition, FP2 – Energy transition and digital transition in urban regeneration and construction and Sapienza Ateneo Research grant “La disintermediazione della Pubblica Amministrazione: il ruolo della tecnologia blockchain e le sue implicazioni nei processi e nei ruoli della PA.” Aris Anagnostopoulos was supported by the ERC Advanced Grant 788893 AMDROMA, the EC H2020RIA project “SoBigData++” (871042), the PNRR MUR project PE0000013-FAIR, the PNRR MUR project IR0000013-SoBigData.it, and the MUR PRIN project 2022EKNE5K “Learning in Markets and Society.” Ioannis Chatzigiannakis was supported by PE07-SERICS (Security and Rights in the Cyberspace) – European Union Next-Generation-EU-PE0000014 (Piano Nazionale di Ripresa e Resilienza – PNRR). Andrea Vitaletti was supported by PE11 - MICS (Made in Italy – Circular and Sustainable) – European Union Next-Generation-EU (Piano Nazionale di Ripresa e Resilienza – PNRR) and the project SERICS (PE00000014) under the MUR National Recovery and Resilience Plan funded by the European Union - NextGenerationEU.

### v2

@@ 10 Design Insights and Opportunities
At the time of writing this work, no analyses or empirical studies about the effect of the spatiotemporal skew on the performance of FL models exist. Thus, for the first time, we produce experiments to understand how different spatiotemporal non-IID data levels affect an FL model’s prediction power. The results show (see Table 12) that high levels of space or time skews decrease the performance of the models, more specifically when the HD is higher than 0.75 (severe degree of non-IID data).
Thus, researchers may benchmark techniques to deal with space and time skew in FL [61, 21, 82] to determine the behavior under high non-IID data levels.
Methods to compare mixed non-IID data types.
@@
Current tools and methods for synthetic partitioning centralized data into federated data [33, 78, 37, 53, 30] focus on simulating one type of non-IID data (label, feature, quantity, spatiotemporal skewness). Nevertheless, a more realistic scenario would be combining two or more types of non-IID data to evaluate the extent to which such mixes can alter the performance of FL models. Therefore, for research in FL purposes, it would be interesting to create methods to partition centralized data into federated clients that permit the control of non-IID data level for two or more data skews simultaneously.
11 Conclusions
@@
This study provides a comprehensive empirical analysis of the non-IID effect in FL. Under controlled conditions, we benchmarked five state-of-the-art strategies for addressing non-IID data distributions, including label, feature, quantity, and spatiotemporal skew, placing particular focus on the relatively unexplored spatiotemporal dimension. We aim to standardize the methodology for studying non-IID data in FL by using HD to quantify data distribution differences. Our findings reveal the significant impact of labels and spatiotemporal skews of non-IID types on FL model performance. We also demonstrate that the model’s performance drop appears at a double threshold. When HD is higher than 0.5 and 0.75, higher damage and a steeper decrease in performance slope occur. Moreover, our results suggest that the FL performance is heavily affected, mainly when the degree of non-IID data is extreme. Thus, we offer valuable recommendations for researchers to address non-IID data. This work represents the most thorough examination of non-IID data in FL to date, providing a robust foundation for future research in FL.
12 Acknowledgments
To be included after acceptance.